OpenStudy (anonymous):

How do I solve: x+4/7=5/6

OpenStudy (jlastino):

is x+4 in the numerator?

OpenStudy (jlastino):

Multiply both sides by the LCD (Which is 42) 42(x+4/7) =(5/6)42 Then divide 42 by the denominators 6(x+4) = 5(7) Distribute 6x+24=35 Then solve all the way 6x=35-24 6x=11 x=11/6

OpenStudy (anonymous):

got it